# MiMo: Unlocking the Reasoning Potential of Language Model – From Pretraining to Posttraining

### v1

@@ 2.1 Pre-Training Data
Web pages naturally contain content with high density reasoning patterns, such as coding tutorial and mathematics blogs. However, we discover that commonly used extractors (Barbaresi, 2021) often fail to preserve mathematics equations and code snippets embedded in the webpage. To address this limitation, we develop a novel HTML-extraction tool specially optimized for mathematics content (Liu et al., 2024b; Paster et al., 2024; Zhou et al., 2025), code blocks, and forum websites. For papers and books, we enhance PDF parsing toolkits to better handle STEM and code content. With these optimized extraction tools, we successfully preserved massive reasoning patterns for subsequent processing stages.
@@ 3.4 RL Infrastructures
We develop the Seamless Rollout Engine and enhance vLLM’s robustness to enable efficient dynamic-sampling-based RL training. We construct our RL system based on verl (Sheng et al., 2024), an open-source RL training library. The library uses Ray (Moritz et al., 2018) to manage computation and communication, implementing the rollout and training phases in Ray Actors and exchanging training data through Ray Objects. Although verl supports flexible implementations of various RL algorithms, it suffers from GPU idle time during both rollout and reward computation phases. Due to the skewness in response lengths, we observe that most GPUs remain idle while waiting for a few long-sequence rollout workers, resulting in wasted computational resources and a slow training process. Several prior works have identified this issue and proposed system-level solutions (Zhong et al., 2024b; Team et al., 2025; Seed et al., 2025). However, most of these solutions rely on asynchronous training, which modifies the underlying algorithm and introduces staleness in long-sequence responses. Rule-based reward computation is also time-consuming, particularly for code data, leading to idle periods for valuable GPU resources. Our use of dynamic sampling, while improving sample efficiency, exacerbates GPU idle time, and leads to wasted examples during multi-turn rollouts. To simultaneously optimize GPU utilization and reduce sample waste, we develop the Seamless Rollout Engine, opportunistically filling sample batches into rollout while performing asynchronous reward computation. Our system builds on the vLLM inference engine (Kwon et al., 2023), and we collaborate with the open-source community to enhance the robustness of vLLM’s “external launch” mode within the verl framework. Additionally, we implement MTP in vLLM to support both MiMo-7B and MiMo-7B-RL.
@@ 4 Conclusion
This work introduces MiMo-7B, a series of LLMs which unlock advanced reasoning capabilities through optimized pre-training and post-training process. Exposed to diverse reasoning patterns during pre-training, MiMo-7B-Base possesses exceptional reasoning potential, outperforming models of significantly larger scale. For post-training, with our robust and efficient RL frameworks, we trained MiMo-7B-RL-Zero and MiMo-7B-RL which demonstrate superior reasoning capabilities across mathematics, code and general tasks. Notably, MiMo-7B-RL achieves 49.3% on LiveCodeBench v6 and 55.4% on AIME 2025, surpassing OpenAI’s o1-mini. We hope this work offers insights for developing more powerful reasoning models.

### v2

@@ 2.1 Pre-Training Data
Web pages naturally contain content with high density reasoning patterns, such as coding tutorial and mathematics blogs. However, we discover that commonly used extractors (Barbaresi, 2021) often fail to preserve mathematics equations and code snippets embedded in the webpage. To address this limitation, we develop a novel HTML-extraction tool specially optimized for mathematics content (Liu et al., 2024c; Paster et al., 2024; Zhou et al., 2025), code blocks, and forum websites. For papers and books, we enhance PDF parsing toolkits to better handle STEM and code content. With these optimized extraction tools, we successfully preserved massive reasoning patterns for subsequent processing stages.
@@ 3.4 RL Infrastructures
We develop the Seamless Rollout Engine and enhance vLLM’s robustness to enable efficient dynamic-sampling-based RL training. We construct our RL system based on verl (Sheng et al., 2024), an open-source RL training library. The library uses Ray (Moritz et al., 2018) to manage computation and communication, implementing the rollout and training phases in Ray Actors and exchanging training data through Ray Objects. Although verl supports flexible implementations of various RL algorithms, it suffers from GPU idle time during both rollout and reward computation phases. Due to the skewness in response lengths, we observe that most GPUs remain idle while waiting for a few long-sequence rollout workers, resulting in wasted computational resources and a slow training process. Several prior works have identified this issue and proposed system-level solutions (Zhong et al., 2024b; Team et al., 2025b; Seed et al., 2025). However, most of these solutions rely on asynchronous training, which modifies the underlying algorithm and introduces staleness in long-sequence responses. Rule-based reward computation is also time-consuming, particularly for code data, leading to idle periods for valuable GPU resources. Our use of dynamic sampling, while improving sample efficiency, exacerbates GPU idle time, and leads to wasted examples during multi-turn rollouts. To simultaneously optimize GPU utilization and reduce sample waste, we develop the Seamless Rollout Engine, opportunistically filling sample batches into rollout while performing asynchronous reward computation. Our system builds on the vLLM inference engine (Kwon et al., 2023), and we collaborate with the open-source community to enhance the robustness of vLLM’s “external launch” mode within the verl framework. Additionally, we implement MTP in vLLM to support both MiMo-7B and MiMo-7B-RL.
@@ 3.6 Discussion
Impact of SFT Data Scaling
Building upon preliminary experiments, our study significantly scaled the SFT dataset from approximately 500K to 6M instances. We empirically observed that this substantial expansion of SFT data resulted in marked improvements in the model’s reasoning abilities and its capacity for generalized dialogue, without compromising its potential for subsequent RL. As detailed in Table 6, the model trained with 6M SFT instances exhibited considerable advancements over its counterpart trained with 500K instances in areas such as mathematical reasoning, code reasoning, scientific reasoning and general dialogue capabilities. Importantly, models subsequently fine-tuned with RL following this enhanced SFT stage also demonstrated sustained performance improvements.
Benchmark MiMo-7B-SFT-500K MiMo-7B-SFT-6M MiMo-7B-RL MiMo-7B-RL-0530
AIME 24 58.7 68.3 68.2 80.1
AIME 25 44.3 50.9 55.4 70.2
MATH500 93.0 94.8 95.8 97.2
GPQA Diamond 50.7 54.1 54.4 60.6
LiveCodeBench v5 52.3 53.4 57.8 60.9
Alignbench v1.1 6.7 7.1 6.9 7.4
Table 6: Model Performance Comparison on Various Benchmarks. MiMo-7B-RL-0530 was evaluated at a 48K context length, its training length, while the other three models were assessed at their 32K training context length. Evaluations for Alignbench v1.1 Liu et al. (2024b) were conducted using GPT-4.1 as the judge.
On-Policy RL with Extended Generation Budget
Our prior empirical investigations indicated that a vanilla implementation of GRPO was markedly prone to premature performance saturation. To mitigate this, we adopted an on-policy RL algorithm, drawing parallels with the approach used in MiMo-VL-7B-RL Team et al. (2025a). Training with on-policy RL proved to be remarkably stable, while also enabling sustained growth in model efficacy throughout the learning process. Further extending our findings, we observed that continuous elevation of the generation length budget during on-policy RL training consistently boosts model performance. Specifically, our RL training protocol involved systematically increasing the model’s generation length from 32K to 38K, and subsequently to 48K. This progressive extension of the generation budget was instrumental in our 7B model ultimately achieving parity with the Deepseek-R1 performance in mathematical reasoning. The MiMo-7B-RL-0530 model has been open-sourced and is publicly available [^1] https://huggingface.co/XiaomiMiMo/MiMo-7B-RL-0530.
Figure 8: MiMo-7B-RL-0530 Performance Curves on AIME24.
@@ 4 Conclusion
This work introduces MiMo-7B, a series of LLMs which unlock advanced reasoning capabilities through optimized pre-training and post-training process. Exposed to diverse reasoning patterns during pre-training, MiMo-7B-Base possesses exceptional reasoning potential, outperforming models of significantly larger scale. For post-training, with our robust and efficient RL frameworks, we trained MiMo-7B-RL-Zero and MiMo-7B-RL which demonstrate superior reasoning capabilities across mathematics, code and general tasks. We hope this work offers insights for developing more powerful reasoning models.